OpenStudy (anonymous):

Find the equation of the line tangent to the graph of the given function at the point with the indicated x-coordinate. f(x)=[((sqr x) + 1)/(sqt x) + 3)]; x=4

OpenStudy (anonymous):

\[f(x)=\frac{ \sqrt{x}+1 }{ \sqrt{x}+3 }; x=4\]

terenzreignz (terenzreignz):

Differentiate the function ^_^

OpenStudy (anonymous):

to find the tangent line?

terenzreignz (terenzreignz):

Look... to find the equation of any line, most of the time, you need a slope and a point. To find the slope of the tangent line, we need the derivative.^_^

terenzreignz (terenzreignz):

Pro-tip... might be a little too late, but meh... If you think doing this integral is a little tedious, use this fact to your advantage: \[\Large \frac{\sqrt x + 1}{\sqrt x + 3}= \frac{\sqrt x + 3 - 2}{\sqrt x + 3}=1 - \frac2{\sqrt x + 3}\]

OpenStudy (anonymous):

I get \[f \prime(x)= 1+\frac{ 3x ^{-\frac{ 1 }{ 2 }} }{ 2(x+6\sqrt{x}+9 }-\frac{ x ^{-\frac{ 1 }{ 2 }}(\sqrt{x}+1) }{ 2(x+6\sqrt{x}+9) }\]

terenzreignz (terenzreignz):

Sorry... I'm not good when it comes to troubleshooting someone else's derivative... (these can get pretty messy at times) Please hold...

terenzreignz (terenzreignz):

Something's off. Try doing it again... you didn't post your solution, so I can't really see where the mistake was. Oh, and don't expand, keep squares as is, they can really reduce headaches ^_^

OpenStudy (anonymous):

Oh I though I could just leave it like that

terenzreignz (terenzreignz):

You could... but you might want to reduce your own (and your instructor's) headache by keeping your notation compact. Anyway, try differentiating instead that one I posted. It's simpler.